• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name brads

    Log entry time 18:05:46 on October 14, 2010

    Entry number 334871

    keyword=Injector matching/tuning was completed before the BCM calibration started

    Trolling through the ELOG, it looks like the injector work was done before the BCM calibration started. This is not the original order of operations that I had been told -- no problem, glad its done, but there seems to be some communication issues here. I'll ask about it tomorrow morning.

    Bottom line is that the Compton background rates should be better, and should not get worse when Hall C goes to high current. I don't think the injector experts understand why the machine is drifting yet though, so the problem will probably return before too long.